I don't understand this problem:
-1/8 + y = -3/4

Found 2 solutions by IWork4Dessert, ewatrrr:
Answer by IWork4Dessert(60)

Just solve:
-1/8 + y = -3/4
+1/8 +1/8
-3/4 + 1/8 = 6/8 - 1/8 = -5/8
y = -5/8
If you didn't get that step above, if you're adding to a negative you're basically decreasing the numerical value. I know that sounds weird, but think of it this way: Whereas with positive numbers 3+5=8, with negative numbers -3+5=2. It's like a ladder that's halfway underwater; even if there are more rungs under the water than above it, all the numbers under the water are less valuable.
Hope this helps!
~IW4D

Answer by ewatrrr(24785)


Hi,
-1/8 + y = -3/4
y = -3/4 + 1/8
y = -6/8 + 1/8
y = -5/8